# Positivstellensätze for polynomial matrices with universal quantifiers

### v2

@@ 6. Proofs
6.1. Proof of Theorem 3.1
Proof of Theorem 3.1.
As mentioned earlier, the proof proceeds in four steps: Step 1: Let $\mathscr{L}$ be the linear functional in Corollary 3.1. The linear functional $\mathscr{L}$ induces a symmetric bilinear form on the space ${\mathbb{R}}[{\boldsymbol{x}}]^{p}$ by
$(\psi,\phi)\mapsto\langle\psi,\phi\rangle_{\mathscr{L}}\coloneqq\frac{1}{2}\mathscr{L}(\psi\phi^{\intercal}+\phi\psi^{\intercal}),\quad\forall\psi,\phi\in{\mathbb{R}}[{\boldsymbol{x}}]^{p}.$
This form is PSD since $\psi\psi^{\intercal}\in\mathcal{Q}^{p}(G,\nu)$ for any $\psi\in{\mathbb{R}}[{\boldsymbol{x}}]^{p}$. Let
$\mathcal{N}\coloneqq\{\psi\in{\mathbb{R}}[{\boldsymbol{x}}]^{p}\mid\langle\psi,\psi\rangle_{\mathscr{L}}=0\}$
be the $\langle\cdot,\cdot\rangle_{\mathscr{L}}$-null vector space corresponding to $\mathscr{L}$. By the Cauchy-Schwarz inequality, $\mathcal{N}$ is a vector subspace of ${\mathbb{R}}[{\boldsymbol{x}}]^{p}$ and $\mathscr{L}$ induces an inner product on the quotient space ${\mathbb{R}}[{\boldsymbol{x}}]^{p}/\mathcal{N}$. Denote by $\mathcal{H}$ the Hilbert space completion of ${\mathbb{R}}[{\boldsymbol{x}}]^{p}/\mathcal{N}$ and $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ its inner product. Recall that by the proof of Lemma 3.1, there exists some $C>0$ such that $C-x_{i}^{2}\in\mathcal{Q}^{1}(G,\nu)$ for all $i\in[n]$. It implies that $(C-x_{i}^{2})\psi\psi^{\intercal}\in\mathcal{Q}^{p}(G,\nu)$ for any $\psi\in{\mathbb{R}}[{\boldsymbol{x}}]^{p}$. Hence,
$0\leq\langle x_{i}\psi+\mathcal{N},x_{i}\psi+\mathcal{N}\rangle_{\mathcal{H}}=\mathscr{L}(x_{i}^{2}\psi\psi^{\intercal})\leq C\mathscr{L}(\psi\psi^{\intercal}).$
Therefore, the multiplication $\hat{x}_{i}:\psi\mapsto x_{i}\psi$ defines a bounded operator on ${\mathbb{R}}[{\boldsymbol{x}}]^{p}/\mathcal{N}$ and can be extended to a bounded operator on $\mathcal{H}$. Fix an orthonormal basis of $\mathcal{H}$ and let $\widehat{A}_{i}$ be the corresponding representative of the multiplication by $x_{i}$ in $\mathcal{B}(\mathcal{H})$ with respect to this basis. Let $\widehat{\mathbf{A}}\coloneqq(\widehat{A}_{i})_{i\in[n]}$.
Denote by $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ the $p\times p$ matrix algebra over the $C^{\star}$-algebra $\mathcal{B}(\mathcal{H})$. That is, $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ consists of $p\times p$ matrices whose entries are in $\mathcal{B}(\mathcal{H})$. Then, $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ is a $C^{\star}$-algebra with involution $[B_{ij}]^{\star}=[B_{ji}^{\star}]$ where $B_{ij}\in\mathcal{B}(\mathcal{H})$. For $k\in[p]$, let ${\boldsymbol{v}}_{k}\in\mathcal{H}$ be the image of the constant polynomial vector $e_{k}\in{\mathbb{R}}^{p}$, where $\{e_{1},\ldots,e_{p}\}$ is the standard basis for ${\mathbb{R}}^{p}$. Define a linear functional $\Pi$ on $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ by
$\Pi(\mathbf{B})=\sum_{1\leq i,j\leq p}\langle B_{ij}{\boldsymbol{v}}_{j},{\boldsymbol{v}}_{i}\rangle_{\mathcal{H}},\ \ \forall\mathbf{B}=[B_{ij}]\in\mathcal{M}_{p}(\mathcal{B}(\mathcal{H})).$
Then, $\Pi$ is a state on $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$. Note that for any $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, it holds that $\Pi(H(\widehat{\mathbf{A}}))=\mathscr{L}(H)$.
Step 2: By applying the GNS construction to $(\mathcal{M}_{p}(\mathcal{B}(\mathcal{H})),\Pi)$, we obtain a Hilbert space $\mathcal{K}$ with an inner product $\langle\cdot,\cdot\rangle_{\mathcal{K}}$, a representation $\pi:\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))\to\mathcal{B}(\mathcal{K})$ and a vector $\xi\in\mathcal{K}$ such that $\Pi(\mathbf{B})=\langle\pi(\mathbf{B})\xi,\xi\rangle_{\mathcal{K}}$ for all $\mathbf{B}\in\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$.
For each $i,j\in[p]$, denote by $E_{ij}$ the $p\times p$ matrix whose $(i,j)$-th entry is the identity operator and all other entries are the zero operator in $\mathcal{B}(\mathcal{H})$. For each $i\in[p]$, let $\mathcal{K}_{i}=\pi(E_{ii})(\mathcal{K})$. Then, the subspaces $\mathcal{K}_{i}$’s are all mutually orthogonal and isomorphic. In fact, $\pi(E_{ij})$ is a linear isomorphism between $\mathcal{K}_{j}$ and $\mathcal{K}_{i}$ with the inverse $\pi(E_{ji})$. Moreover,
$\mathcal{K}=1_{\mathcal{B}(\mathcal{K})}\mathcal{K}=\pi\left(\sum_{i=1}^{p}E_{ii}\right)\mathcal{K}=\bigoplus_{i=1}^{p}\mathcal{K}_{i}.$
Hence, each $\mathcal{K}_{i}$ is a Hilbert subspace of $\mathcal{K}$. Let us decompose $\xi$ as $\xi=\xi_{1}+\cdots+\xi_{p}$ where each $\xi_{i}\in\mathcal{K}_{i}$, i.e., $\xi_{i}=\pi(E_{ii})\xi$.
For $H({\boldsymbol{x}})=[H_{ij}({\boldsymbol{x}})]\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, it holds that
$\mathscr{L}(H)$ $=\Pi(H(\widehat{\mathbf{A}}))=\left\langle\pi\left(\sum_{i,j}H_{ij}(\widehat{\mathbf{A}})E_{ij}\right)\sum_{\ell=1}^{p}\xi_{\ell},\sum_{\ell=1}^{p}\xi_{\ell}\right\rangle_{\mathcal{K}}$ (9)
$=\left\langle\sum_{i,j}\left(\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\pi(E_{ij})\right)\sum_{\ell=1}^{p}\xi_{\ell},\sum_{\ell=1}^{p}\xi_{\ell}\right\rangle_{\mathcal{K}}=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\pi(E_{ij})\xi_{j},\xi_{i}\right\rangle_{\mathcal{K}}.$
For $i,j\in[p]$, let $\eta_{j}=\pi(E_{1i})\pi(E_{ij})\xi_{j}$ and $\eta_{i}=\pi(E_{1i})\xi_{i}$. Then, $\eta_{j},\eta_{i}\in\mathcal{K}_{1}$ and
$\mathscr{L}(H)$ $=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\pi(E_{i1})\eta_{j},\pi(E_{i1})\eta_{i}\right\rangle_{\mathcal{K}}=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\eta_{j},\pi(E_{1i})\pi(E_{i1})\eta_{i}\right\rangle_{\mathcal{K}}$
$=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\eta_{j},\pi(E_{11})\eta_{i}\right\rangle_{\mathcal{K}}=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\eta_{j},\eta_{i}\right\rangle_{\mathcal{K}}.$
For each $i\in[p]$, let $\mathbf{A}=(A_{1},\ldots,A_{n})$ with $A_{i}\in\mathcal{B}(\mathcal{K}_{1})$ being the restriction of $\pi(\widehat{A}_{i}I_{p})$ to $\mathcal{K}_{1}$ and denote by $\langle\cdot,\cdot\rangle_{\mathcal{K}_{1}}$ the inner product on $\mathcal{K}_{1}$ induced from $\langle\cdot,\cdot\rangle_{\mathcal{K}}$. Then, we have
$\mathscr{L}(H)=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\eta_{j},\eta_{i}\right\rangle_{\mathcal{K}}=\sum_{i,j}\left\langle H_{ij}(\mathbf{A})\eta_{j},\eta_{i}\right\rangle_{\mathcal{K}_{1}}.$ (10)
Step 3: For each $i\in[n]$, $A_{i}\in\mathcal{B}(\mathcal{K}_{1})$ is self-adjoint. Indeed, for every $\omega,\gamma\in\mathcal{K}_{1}$,
$\langle A_{i}\omega,\gamma\rangle_{\mathcal{K}_{1}}=\langle\pi(\widehat{A}_{i}I_{p})\omega,\gamma\rangle_{\mathcal{K}}=\langle\omega,(\pi(\widehat{A}_{i}I_{p}))^{\star}\gamma\rangle_{\mathcal{K}}=\langle\omega,\pi((\widehat{A}_{i}I_{p})^{\star})\gamma\rangle_{\mathcal{K}}=\langle\omega,\pi(\widehat{A}_{i}I_{p})\gamma\rangle_{\mathcal{K}}=\langle\omega,A_{i}\gamma\rangle_{\mathcal{K}_{1}}.$
Moreover, $(A_{i})_{i\in[n]}$ are pairwise commute. In fact, for every $\omega,\gamma\in\mathcal{K}_{1}$,
$\langle A_{i}A_{j}\omega,\gamma\rangle_{\mathcal{K}_{1}}$ $=\langle\pi(\widehat{A}_{i}I_{p})\pi(\widehat{A}_{j}I_{p})\omega,\gamma)\rangle_{\mathcal{K}}=\langle\pi(\widehat{A}_{i}\widehat{A}_{j}I_{p})\omega,\gamma)\rangle_{\mathcal{K}}$
$=\langle\pi(\widehat{A}_{j}\widehat{A}_{i}I_{p})\omega,\gamma)\rangle_{\mathcal{K}}=\langle\pi(\widehat{A}_{j}I_{p})\pi(\widehat{A}_{i}I_{p})\omega,\gamma)\rangle_{\mathcal{K}}=\langle A_{j}A_{i}\omega,\gamma\rangle_{\mathcal{K}_{1}}.$
For each $i\in[n]$, as $A_{i}\in\mathcal{B}(\mathcal{K}_{1})$ is self-adjoint, by the spectral theorem for bounded self-adjoint operators on a Hilbert space (see, e.g. [48, Theorem 5.1]), there exists a unique projection-valued measure $\mathcal{E}_{i}$ on the Borel $\sigma$-algebra $\mathfrak{B}({\mathbb{R}})$ such that the support of $\mathcal{E}_{i}$ is contained in a compact interval, $A_{i}=\int_{{\mathbb{R}}}t\mathrm{d}\mathcal{E}_{i}(t)$ and
$\langle A_{i}\omega,\gamma\rangle_{\mathcal{K}_{1}}=\int_{{\mathbb{R}}}t\mathrm{d}\langle\mathcal{E}_{i}(t)\omega,\gamma\rangle_{\mathcal{K}_{1}},\quad\forall\omega,\gamma\in\mathcal{K}_{1}.$
Moreover, as $(A_{i})_{i\in[n]}$ are pairwise commute, by [48, Theorem 4.10], we can define on the Borel $\sigma$-algebra $\mathfrak{B}({\mathbb{R}}^{n})$ the product measure $\mathcal{E}$ of $\mathcal{E}_{i},i\in[n]$, such that $\mathcal{E}\left(\prod_{i\in[n]}\mathcal{N}_{i}\right)=\prod_{i\in[n]}\mathcal{E}_{i}(\mathcal{N}_{i})$ for all $\mathcal{N}_{i}\in\mathfrak{B}({\mathbb{R}}),\ i\in[n].$ By the multidimensional spectral theorem (see, [48, Theorem 5.23] or [4, Sec. 6.5, Theorem 1]), $A_{i}=\int_{{\mathbb{R}}^{n}}x_{i}\mathrm{d}\mathcal{E}({\boldsymbol{x}})$, $i\in[n]$, and
$\left\langle\left(\prod_{i\in[n]}A_{i}^{\alpha_{i}}\right)\omega,\gamma\right\rangle_{\mathcal{K}_{1}}=\int_{{\mathbb{R}}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\langle\mathcal{E}({\boldsymbol{x}})\omega,\gamma\rangle_{\mathcal{K}_{1}},\quad\forall{\boldsymbol{\alpha}}\in\mathbb{N}^{n},\ \omega,\gamma\in\mathcal{K}_{1}.$
Note that the support of the product measure $\mathcal{E}$ is compact since each $\mathcal{E}_{i}$ has compact support.
For $i,j\in[p]$, recall that the vectors $\eta_{i},\eta_{j}\in\mathcal{K}_{1}$. Define the real-valued measure $\phi_{ij}(\cdot)\coloneqq\langle\mathcal{E}(\cdot)\eta_{j},\eta_{i}\rangle_{\mathcal{K}_{1}}$ on the Borel $\sigma$-algebra $\mathfrak{B}({\mathbb{R}}^{n})$ and the symmetric matrix-valued measure $\Phi\coloneqq[\phi_{ij}]$. Then, each $\phi_{ij}$ has a compact support and for any $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, we have $\mathscr{L}(H)=\int_{{\mathbb{R}}^{n}}H({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})$ by (10). Moreover, $\Phi$ is a PSD matrix-valued measure. In fact, for any $\mathcal{N}\in\mathfrak{B}({\mathbb{R}}^{n})$ and ${\boldsymbol{v}}=(v_{i})\in{\mathbb{R}}^{p}$, as $\mathcal{E}(\mathcal{N})$ is a projection operator on $\mathcal{K}_{1}$, we have
${\boldsymbol{v}}^{\intercal}\Phi(\mathcal{N}){\boldsymbol{v}}=\sum_{i,j}\phi_{ij}(\mathcal{N})v_{i}v_{j}=\sum_{i,j}\langle\mathcal{E}(\mathcal{N})v_{j}\eta_{j},v_{i}\eta_{i}\rangle_{\mathcal{K}_{1}}=\left\langle\mathcal{E}(\mathcal{N})\sum_{i}v_{i}\eta_{i},\mathcal{E}(\mathcal{N})\sum_{j}v_{j}\eta_{j}\right\rangle_{\mathcal{K}_{1}}\geq 0.$
Step 4: Now we show that for $i,j\in[p]$, the support of $\phi_{ij}$ is contained in $\mathcal{X}$. Suppose that ${\boldsymbol{u}}\in{\mathbb{R}}^{n}$ and ${\boldsymbol{u}}\not\in\mathcal{X}$, i.e., there exists ${\boldsymbol{w}}\in\mathcal{Y}$ such that $G({\boldsymbol{u}},{\boldsymbol{w}})\not\succeq 0$. Then, there is a ball $\mathcal{O}_{1}\subset{\mathbb{R}}^{n}$ (resp., $\mathcal{O}_{2}\subset{\mathbb{R}}^{m}$) with a radius $\rho>0$ around ${\boldsymbol{u}}$ (resp., ${\boldsymbol{w}}$) such that $G({\boldsymbol{x}},{\boldsymbol{y}})\not\succeq 0$ on $\mathcal{O}_{1}\times\mathcal{O}_{2}$. Without loss of genericity, we may assume that there exists ${\boldsymbol{v}}\in{\mathbb{R}}^{q}$ and $\delta>0$ such that ${\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}\leq-\delta$ on $2\mathcal{O}_{1}\times 2\mathcal{O}_{2}$. Define a continuous function $h^{(1)}({\boldsymbol{x}})$ (resp. $h^{(2)}({\boldsymbol{y}})$) on ${\mathbb{R}}^{n}$ (resp., ${\mathbb{R}}^{m}$) by $h^{(1)}({\boldsymbol{x}})=2\rho-\|{\boldsymbol{x}}-{\boldsymbol{u}}\|$ (resp., $h^{(2)}({\boldsymbol{y}})=2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|$)for ${\boldsymbol{x}}\in 2\mathcal{O}_{1}$ (resp., ${\boldsymbol{y}}\in 2\mathcal{O}_{2}$) and $h^{(1)}({\boldsymbol{x}})=0$ (resp., $h^{(2)}({\boldsymbol{y}})=0$) otherwise. Take a compact set $\mathcal{C}\subset{\mathbb{R}}^{n}$ containing $2\mathcal{O}_{1}$ and the support of $\phi_{ii}$, $i\in[p]$. By the Weierstrass theorem, there is a sequence of square polynomials $\{h^{2}_{\ell}\}_{\ell\in\mathbb{N}}\subset{\mathbb{R}}[{\boldsymbol{x}}]$ which converges to $h^{(1)}$ uniformly on $\mathcal{C}$. By Proposition 3.1, there exists a sequence of SOS polynomilas $\{\sigma_{k}\}_{k}$ in ${\mathbb{R}}[{\boldsymbol{y}}]$ that converges to $h^{(2)}$ in the $L^{2}$-norm. Hence,
$\lim_{\ell\to\infty}\lim_{k\to\infty}\sum_{i=1}^{p}\int_{{\mathbb{R}}^{n}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$ (11)
$=$ $\lim_{\ell\to\infty}\lim_{k\to\infty}\sum_{i=1}^{p}\int_{\mathcal{C}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
$=$ $\lim_{\ell\to\infty}\sum_{i=1}^{p}\int_{\mathcal{C}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h_{\ell}^{2}({\boldsymbol{x}})h^{(2)}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
$=$ $\sum_{i=1}^{p}\int_{\mathcal{C}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h^{(1)}({\boldsymbol{x}})h^{(2)}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
$=$ $\sum_{i=1}^{p}\int_{2\mathcal{O}_{1}}\left(\int_{\mathcal{Y}\cap 2\mathcal{O}_{2}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}(2\rho-\|{\boldsymbol{x}}-{\boldsymbol{u}}\|)(2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|)\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
$\leq$ $\sum_{i=1}^{p}\int_{2\mathcal{O}_{1}}\left(\int_{\mathcal{Y}\cap\mathcal{O}_{2}}-\delta\rho(2\rho-\|{\boldsymbol{x}}-{\boldsymbol{u}}\|)\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
$=$ $-\delta\rho\nu(\mathcal{Y}\cap\mathcal{O}_{2})\sum_{i=1}^{p}\int_{2\mathcal{O}_{1}}(2\rho-\|{\boldsymbol{x}}-{\boldsymbol{u}}\|)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
$\leq$ $-\delta\rho\nu(\mathcal{Y}\cap\mathcal{O}_{2})\sum_{i=1}^{p}\int_{\mathcal{O}_{1}}\rho\mathrm{d}\phi_{ii}({\boldsymbol{x}})=-\delta\rho^{2}\nu(\mathcal{Y}\cap\mathcal{O}_{2})\sum_{i=1}^{p}\phi_{ii}(\mathcal{O}_{1}).$
As $\mathcal{C}$ is compact, the second and third equalities in (11) can be verified by Dominated Convergence Theorem and the convergence of $\{\sigma_{k}\}_{k}$ to $h^{(2)}$ in the $L^{2}$-norm.
Note that for all $\ell,k\in\mathbb{N}$,
$\sum_{i=1}^{p}\int_{{\mathbb{R}}^{n}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$ (12)
$=$ $\int_{{\mathbb{R}}^{n}}\left(\int_{\mathcal{Y}}\left\langle h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\Phi({\boldsymbol{x}})$
$=$ $\mathscr{L}\left(\int_{\mathcal{Y}}\left\langle h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\geq 0,$
where the last inequality is due to the fact that $h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}$ is an SOS and $\mathscr{L}(\mathcal{Q}^{p}(G,\nu))\subseteq{\mathbb{R}}_{+}$. Since $\mathrm{supp}(\nu)=\mathcal{Y}$, we have $\nu(\mathcal{Y}\cap\mathcal{O}_{2})>0$. Therefore, by combining (11) and (12), we have $\phi_{ii}(\mathcal{O}_{1})=0$ for all $i\in[p]$. As $\Phi$ is PSD, it implies that $\phi_{ij}(\mathcal{O}_{1})=0$ for all $i,j\in[p]$. Hence, the support of each $\phi_{ij}$ is contained in $\mathcal{X}$. Then, as $F$ is positive on $\mathcal{X}$ by assumption and $\Phi$ is PSD, we have
$0\geq\mathscr{L}(F)=\int_{\mathcal{X}}F({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})>0,$
a contradiction. ∎
@@
$M\coloneqq\max\left\{|\lambda_{\max}(G_{j}({\boldsymbol{x}},{\boldsymbol{y}}))|\colon{\boldsymbol{x}}\in[-C,C]^{n},\ {\boldsymbol{y}}\in\mathcal{Y},\ j\in[s]\right\}.$
As $\mathcal{Y}$ is compact, the quantity $M$ is well-defined. For this $M$, we show that there exists some $\bar{k}\in\mathbb{N}$ such that (4) holds for all $k\geq\bar{k}$. Suppose on the contrary that for any $k\in\mathbb{N}$, there exists ${\boldsymbol{x}}^{(k)}\in[-C,C]^{n}$ such that
@@
As $[-C,C]^{n}$ is compact, without loss of generality, we may assume that $\lim_{k\to\infty}{\boldsymbol{x}}^{(k)}={\boldsymbol{x}}^{\star}$ for some ${\boldsymbol{x}}^{\star}\in[-C,C]^{n}$. Next, we prove that there exists $k^{\prime}\in\mathbb{N}$ and a neighborhood $\mathcal{O}_{1}$ of ${\boldsymbol{x}}^{\star}$ such that (4) holds for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$, which yields a contradiction.
We first consider the case that ${\boldsymbol{x}}^{\star}\in\widehat{\mathcal{X}}$. As $F({\boldsymbol{x}})\succ 0$ on $\widehat{\mathcal{X}}$, there exists $\varepsilon>0$ and a neighborhood $\mathcal{O}_{1}$ of ${\boldsymbol{x}}^{\star}$ such that $F({\boldsymbol{x}})\succeq\varepsilon I_{p}$ for all ${\boldsymbol{x}}\in\mathcal{O}_{1}$. We now prove that there exsits $k^{\prime}\in\mathbb{N}$ such that for all $k\geq k^{\prime}$, it holds
@@
[…]
@@
$\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\leq M\sum_{\lambda_{i}^{(j)}>0}\frac{\lambda_{i}^{(j)}}{M}\left(1-\frac{\lambda_{i}^{(j)}}{M}\right)^{2k}\leq Mq_{j}\frac{\varepsilon}{2q^{\prime}sM\nu(\mathcal{Y})}\leq\frac{\varepsilon}{2s\nu(\mathcal{Y})}$ (15)
@@
Next we consider the case that ${\boldsymbol{x}}^{\star}\in[-C,C]^{n}\setminus\widehat{\mathcal{X}}$. There exists $j_{0}\in[s]$ and ${\boldsymbol{y}}^{(0)}\in\mathcal{Y}$ such that $\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}}^{\star},{\boldsymbol{y}}^{(0)})<0$. By continuity, for some $\bar{\lambda}<0$, there exists a neighborhood $\mathcal{O}_{1}$ (resp., $\mathcal{O}_{2}$) of ${\boldsymbol{x}}^{\star}$ (resp., ${\boldsymbol{y}}^{(0)}$) such that $\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}}))\leq\bar{\lambda}$ for all ${\boldsymbol{x}}\in\mathcal{O}_{1}$ and ${\boldsymbol{y}}\in\mathcal{O}_{2}$. For any pair of ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and ${\boldsymbol{y}}\in\mathcal{O}_{2}\cap\mathcal{Y}$, letting $\lambda_{\min}^{(j_{0})}\coloneqq\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}}))$, from (14) we obtain that
$\left\langle\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle$ $\leq\sum_{\lambda_{i}^{(j_{0})}>0}\lambda_{i}^{(j_{0})}+\sum_{\lambda_{i}^{(j_{0})}<0}\lambda_{i}^{(j_{0})}\left(1-\frac{\lambda_{i}^{(j_{0})}}{M}\right)^{2k}$

### v4

@@ 6. Proofs
$M\coloneqq\max\left\{|\lambda_{\max}(G_{j}({\boldsymbol{x}},{\boldsymbol{y}}))|,|\lambda_{\min}(G_{j}({\boldsymbol{x}},{\boldsymbol{y}}))|\colon{\boldsymbol{x}}\in[-C,C]^{n},\ {\boldsymbol{y}}\in\mathcal{Y},\ j\in[s]\right\}.$
As $\mathcal{Y}$ is compact, the quantity $M$ is well-defined and positive. For this $M$, we show that there exists some $\bar{k}\in\mathbb{N}$ such that (5) holds for all $k\geq\bar{k}$. Suppose on the contrary that for any $k\in\mathbb{N}$, there exists ${\boldsymbol{x}}^{(k)}\in[-C,C]^{n}$ such that
@@
As $C,C]^{n}$ is compact, without loss of generality, we may assume that $\lim_{k\to\infty}{\boldsymbol{x}}^{(k)}={\boldsymbol{x}}^{\star}$ for some ${\boldsymbol{x}}^{\star}\in[-C,C]^{n}$. Next, we claim that there exists $k^{\prime}\in\mathbb{N}$ and a neighborhood $\mathcal{O}_{1}$ of $\boldsymbol{x}$$^{\star}$ such that (5) holds for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$, which yields a contradiction.
We first consider the case that $^{\boldsymbol{x}}^{\star}\in\widehat{\mathcal{X}}$. As $F({\boldsymbol{x}})\succ 0$ on $\widehat{\mathcal{X}}$, there exists $\varepsilon>0$ and a neighborhood $\mathcal{O}_{1}$ of $\boldsymbol{x}$$^{\star}$ such that $F({\boldsymbol{x}})\succeq\varepsilon I_{p}$ for all ${\boldsymbol{x}}\in\mathcal{O}_{1}$. We now prove that there exsits $k^{\prime}\in\mathbb{N}$ such that for all $k\geq k^{\prime}$, it holds
@@
$\begin{aligned} \left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{u}},{\boldsymbol{w}})/M\right)^{2k},G_{j}({\boldsymbol{u}},{\boldsymbol{w}})\right\rangle=&\left\langle\left(I_{q_{j}}-Q_{j}(D_{j}/M)Q_{j}^{\intercal}\right)^{2k},Q_{j}D_{j}Q_{j}^{\intercal}\right\rangle\\ =&\left\langle Q_{j}\left(I_{q_{j}}-D_{j}/M\right)^{2k}Q_{j}^{\intercal},Q_{j}D_{j}Q_{j}^{\intercal}\right\rangle=M\sum_{i=1}^{q_{j}}\frac{\lambda_{i}^{(j)}}{M}\left(1-\frac{\lambda_{i}^{(j)}}{M}\right)^{2k}\end{aligned}.$ (18)
@@
$\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\leq Mq_{j}\frac{\varepsilon}{2q^{\prime}sM\nu(\mathcal{Y})}\leq\frac{\varepsilon}{2s\nu(\mathcal{Y})}$ (19)
@@
Next, we consider the case that $^{\boldsymbol{x}}^{\star}\in[-C,C]^{n}\setminus\widehat{\mathcal{X}}$. There exists $j_{0}\in[s]$ and ${\boldsymbol{y}}^{(0)}\in\mathcal{Y}$ such that $\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}}^{\star},{\boldsymbol{y}}^{(0)})<0$. By continuity, for some $\bar{\lambda}<0$, there exists a neighborhood $\mathcal{O}_{1}$ (resp., $\mathcal{O}_{2}$) of $\boldsymbol{x}$$^{\star}$ (resp., ${\boldsymbol{y}}^{(0)}$) such that $\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}}))\leq\bar{\lambda}$ for all ${\boldsymbol{x}}\in\mathcal{O}_{1}$ and ${\boldsymbol{y}}\in\mathcal{O}_{2}$. For any pair of ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and ${\boldsymbol{y}}\in\mathcal{O}_{2}\cap\mathcal{Y}$, letting $\lambda_{\min}^{(j_{0})}\coloneqq\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}}))$, from (18) we obtain that
$\left\langle\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle$ $\leq\sum_{\lambda_{i}^{(j_{0})}>0}\lambda_{i}^{(j_{0})}+\sum_{\lambda_{i}^{(j_{0})}<0}\lambda_{i}^{(j_{0})}\left(1-\frac{\lambda_{i}^{(j_{0})}}{M}\right)^{2k}$ (20)